How much 5% acid must be mixed with 25% acid to obtain 500 ml of 7% acid
Answer by stanbon(75887)

How much 5% acid must be mixed with 25% acid to obtain 500 ml of 7% acid
------
Equation:
acid + acid = acid
-----
0.05x + 0.25(500-x) = 0.07*500
------
Multiply thru by 100 to get:
5x + 25*500 - 25x = 7*500
-20x = -18*500
x = 18*25
x = 450 ml (amt of 5% solution needed)
---
500 -x = 50 ml (amt. of 25% solution needed)
====================
Cheers,
Stan H.
====================